Calculate the heat needed at 0∘C∘to make each of the following changes of state.

Calories to melt 22 g of ice
1
Expert's answer
2020-09-14T08:29:11-0400

QUESTION # 132550

Calculate the heat needed at 0∘C∘to make each of the following changes of state.

Calories to melt 22 g of ice

ANSWERS

ΔHf = q/m

Where;

ΔHf = enthalpy of fusion of water(80cal/gram)

q =heat needed to melt ice

m = mass of ice

ΔHf = q/m

80cal/g =q/22g

 q = 22g×80cal/g

   =1760 calories